# Dynamic Prompt Generation for Interactive 3D Medical Image Segmentation Training

### v1

@@ 1.2 Related Work
Foundation models for image segmentation, particularly SAM [5] and SAM2 [11], have demonstrated strong zero-shot performance on natural images. Medical adaptations including MedSAM [7] and MedSAM2 [9] fine-tune these models on medical datasets but exhibit limited interactive refinement capabilities, particularly for iterative correction workflows.
Recent advances in interactive 3D medical segmentation include SegVol [1], SAM-Med3D [13], VISTA3D [3], and nnInteractive [2]. These methods explore different approaches to incorporating user feedback in volumetric medical image analysis, establishing the foundation for interactive refinement in clinical workflows.
@@
Existing approaches exhibit fundamental limitations: VISTA3D [3] supports only click prompts, lacking bounding box and iterative refinement capabilities. SegVol [1] accepts only bounding boxes without click-based correction. SAM-Med3D [13] operates at low resolution, insufficient for accurate 3D segmentation. nnInteractive [2] accepts multiple prompts but restricts bounding boxes to 2D slices, ignoring 3D spatial context.
@@ 2.1 General Network Architecture
We base our solution on the nnInteractive [2] method and adopt a 3D Residual Encoder U-Net [12] architecture, combining the spatial precision of U-Net with the gradient flow benefits of residual connections, specifically optimized for interactive volumetric medical image segmentation.
@@ 2.2 Prompt Encoding
Following the nnInteractive [2] strategy, we encode heterogeneous prompts (bounding boxes, positive/negative clicks and previous segmentation masks) as additional input channels concatenated with the image data, as illustrated in Figure 2.
@@ 2.5.3 Inference Optimization
We followed the Docker optimizations in DAFT [10] and reduced the size and number of layers in the image by: (i) using a minimal base image (python : 3.13-slim-bookworm), (ii) disabling pip caching (PIP_NO_CACHE_DIR=1), and (iii) squashing layers with docker-squash.
@@ 2.6 Loss Function
We employ a compound loss function combining Dice loss and cross-entropy, as such hybrid approaches have proven robust across various medical image segmentation tasks [6]. The total loss is formulated as:
@@ 3.1 Dataset and evaluation metrics
The development set is an extension of the CVPR 2024 MedSAM on Laptop Challenge [8], including more 3D cases from public datasets [^1] A complete list is available at https://medsam-datasetlist.github.io/ and covering commonly used 3D modalities, such as Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Positron Emission Tomography (PET), Ultrasound, and Microscopy images. The hidden testing set is created by a community effort where all the cases are unpublished. The annotations are either provided by the data contributors or annotated by the challenge organizer with 3D Slicer [4] and MedSAM2 [9]. In addition to using all training cases, the challenge contains a coreset track, where participants can select 10% of the total training cases for model development.
@@ 3.2.1 Preprocessing
Following the practice in MedSAM [7], all images were processed to npz format with an intensity range of $[0,255]$. Specifically, for CT images, we initially normalized the Hounsfield units using typical window width and level values: soft tissues (W:400, L:40), lung (W:1500, L:-160), brain (W:80, L:40), and bone (W:1800, L:400). Subsequently, the intensity values were rescaled to the range of $[0,255]$. For other images, we clipped the intensity values to the range between the 0.5th and 99.5th percentiles before rescaling them to the range of [0, 255]. If the original intensity range is already in $[0,255]$, no preprocessing was applied.
@@ 5.0.1 Acknowledgements
We thank all the data owners for making the medical images publicly available and CodaLab [14] for hosting the challenge platform.

### v2

@@ 1.2 Related Work
Foundation models for image segmentation, particularly SAM [SAM-ICCV23] and SAM2 [SAM2-2025-ICLR], have demonstrated strong zero-shot performance on natural images. Medical adaptations including MedSAM [MedSAM] and MedSAM2 [MedSAM2] fine-tune these models on medical datasets but exhibit limited interactive refinement capabilities, particularly for iterative correction workflows.
Recent advances in interactive 3D medical segmentation include SegVol [segvol], SAM-Med3D [SAM-Med3D], VISTA3D [VISTA3D], and nnInteractive [nnInteractive]. These methods explore different approaches to incorporating user feedback in volumetric medical image analysis, establishing the foundation for interactive refinement in clinical workflows.
@@
Existing approaches exhibit fundamental limitations: VISTA3D [VISTA3D] supports only click prompts, lacking bounding box and iterative refinement capabilities. SegVol [segvol] accepts only bounding boxes without click-based correction. SAM-Med3D [SAM-Med3D] operates at low resolution, insufficient for accurate 3D segmentation. nnInteractive [nnInteractive] accepts multiple prompts but restricts bounding boxes to 2D slices, ignoring 3D spatial context.
@@ 2.1 General Network Architecture
We base our solution on the nnInteractive [nnInteractive] method and adopt a 3D Residual Encoder U-Net [U-Net] architecture, combining the spatial precision of U-Net with the gradient flow benefits of residual connections, specifically optimized for interactive volumetric medical image segmentation.
@@ 2.2 Prompt Encoding
Following the nnInteractive [nnInteractive] strategy, we encode heterogeneous prompts (bounding boxes, positive/negative clicks and previous segmentation masks) as additional input channels concatenated with the image data, as illustrated in Figure 2.
@@ 2.5.3 Inference Optimization
We followed the Docker optimizations in DAFT [pfefferle2025daft] and reduced the size and number of layers in the image by: (i) using a minimal base image (python : 3.13-slim-bookworm), (ii) disabling pip caching (PIP_NO_CACHE_DIR=1), and (iii) squashing layers with docker-squash.
@@ 2.6 Loss Function
We employ a compound loss function combining Dice loss and cross-entropy, as such hybrid approaches have proven robust across various medical image segmentation tasks [LossOdyssey]. The total loss is formulated as:
@@ 3.1 Dataset and evaluation metrics
The development set is an extension of the CVPR 2024 MedSAM on Laptop Challenge [CVPR24-EffMedSAMs-summary], including more 3D cases from public datasets [^1] A complete list is available at https://medsam-datasetlist.github.io/ and covering commonly used 3D modalities, such as Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Positron Emission Tomography (PET), Ultrasound, and Microscopy images. The hidden testing set is created by a community effort where all the cases are unpublished. The annotations are either provided by the data contributors or annotated by the challenge organizer with 3D Slicer [slicer] and MedSAM2 [MedSAM2]. In addition to using all training cases, the challenge contains a coreset track, where participants can select 10% of the total training cases for model development.
@@ 3.2.1 Preprocessing
Following the practice in MedSAM [MedSAM], all images were processed to npz format with an intensity range of $[0,255]$. Specifically, for CT images, we initially normalized the Hounsfield units using typical window width and level values: soft tissues (W:400, L:40), lung (W:1500, L:-160), brain (W:80, L:40), and bone (W:1800, L:400). Subsequently, the intensity values were rescaled to the range of $[0,255]$. For other images, we clipped the intensity values to the range between the 0.5th and 99.5th percentiles before rescaling them to the range of [0, 255]. If the original intensity range is already in $[0,255]$, no preprocessing was applied.
@@ 5.0.1 Acknowledgements
We thank all the data owners for making the medical images publicly available and CodaLab [codabench] for hosting the challenge platform.